2. Patients arrive at a hospital accident and emergency department at random follow a Poisson distribution at a rate of 6 per hour.
(a) Find the probability that, during any 90-minute period, the number of patients arriving at the hospital accident and emergency department is
(i) exactly 7 (2 marks) (ii) at least 3 (2 marks)
(b) What is the mean and standard deviation of the number of patients arriving at the hospital accident and emergency department in an hour? (2 marks)
1
Expert's answer
2020-11-24T17:37:28-0500

Let "X=" the number of patients: "X\sim Po(\lambda t)"


"P(X=x)=\dfrac{e^{-\lambda t}(\lambda t)^x}{x!}"

(a) Given "\lambda=6, t=1.5"

(i)


"P(X=7)=\dfrac{e^{-9}(9)^7}{7!}\approx0.11712"

(ii)


"P(X\geq3)=1-P(X=0)-P(X=1)-P(X=2)="

"=1-\dfrac{e^{-9}(9)^0}{0!}-\dfrac{e^{-9}(9)^1}{1!}-\dfrac{e^{-9}(9)^2}{2!}="

"=1-50.5e^{-9}\approx0.9937678"

(b)


"mean=variance=\lambda=6"

"standard \ deviation=\sqrt{\lambda}=\sqrt{6}\approx2.44949"